40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Mar 31 02:51:39 2015, ericq, Update, LSC, Some locks 
    Reply  Tue Mar 31 03:02:44 2015, Koji, Update, LSC, Some locks 
       Reply  Tue Mar 31 18:27:58 2015, ericq, Update, LSC, Some locks qpdSwitch.pngqpdAlways.png
Message ID: 11182     Entry time: Tue Mar 31 02:51:39 2015     Reply to this: 11183
Author: ericq 
Type: Update 
Category: LSC 
Subject: Some locks 

I had a handful of ~10 minute locks tonight. I intended to work on the 1f PRMI transition, but ended just familiarizing myself with the current scheme. 

Before touching anything, I committed the locking scripts to the svn. Unfortunately, the up script as I found it never worked for me tonight. I had to reintroduce the digitial crossover helper in CM_SLOW to get past the ramping up of the overall REFL11 gain. (With this is in place, there is some bad ringing around 200Hz for a time, but it goes away... or unlocks)

I did phase the PD formally known as REFL55 with an 800Hz PRM excitation while in full lock.42 to 102 degrees, ~30dB ratio between the I and Q peaks. However, come to think of it, how much does the CARM loop interfere with this?

The locklosses I had seemed to be due to a large fluctuation in all cavities' power. Maybe this will be helped by better PRC angular control, but we could maybe be helped by normalizing the digital part of the CARM loop by the arm transmissions once lock is acquired. 

ELOG V3.1.3-